# Optimal Control of Nonlinear Systems with Unknown Dynamics

### v2

@@ 3.3 Analysis
In this subsection, we conduct a convergence analysis of the proposed framework with respect to the policy parameter $\boldsymbol{\theta}^{\mu}$ based on the given dataset $\mathcal{D}$. To proceed, we formally define the optimal solutions that this study seeks to determine, as presented in the following definition:
Definition 1 (Optimal solutions)
Parameters $\boldsymbol{\theta}^{f*}\in\mathbb{R}^{p}$, $\boldsymbol{\theta}^{J*}\in\mathbb{R}^{s}$, and $\boldsymbol{\theta}^{\mu*}\in\mathbb{R}^{q}$ are optimal solutions if ${\mathbf{L}}_{f}(A^{*},B^{*},C^{*},\boldsymbol{\theta}^{f*})\leq{\mathbf{L}}_{f}(A,B,C,\boldsymbol{\theta}^{f})$, ${\mathbf{L}}_{J}(\boldsymbol{\theta}^{J*})\leq{\mathbf{L}}_{J}(\boldsymbol{\theta}^{J})$, $\hat{J}_{\mu,t^{*}}(\boldsymbol{\theta}^{f*},\boldsymbol{\theta}^{J*},\boldsymbol{\theta}^{\mu*})\leq\hat{J}_{\mu,t^{*}}(\boldsymbol{\theta}^{f*},\boldsymbol{\theta}^{J*},\boldsymbol{\theta}^{\mu})$, $\nabla_{\boldsymbol{\theta}^{f}}{\mathbf{L}}_{f}(A^{*},B^{*},C^{*},\boldsymbol{\theta}^{f*})=0$, $\nabla_{\boldsymbol{\theta}^{J}}{\mathbf{L}}_{J}(\boldsymbol{\theta}^{J*})=0$, $\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu,t^{*}}(\boldsymbol{\theta}^{f*},\boldsymbol{\theta}^{J*},\boldsymbol{\theta}^{\mu*})=0$, where $A^{*},B^{*}$ and $C^{*}$ are constant matrices assumed to have full row rank, computed according to (13) with $\boldsymbol{\theta}_{k}^{f}$ replaced by $\boldsymbol{\theta}^{f*}$.
For analytical tractability, we impose the following technical assumption:
@@
The functions $\boldsymbol{g}(\boldsymbol{x},\boldsymbol{\theta}^{f})$ and $V(\boldsymbol{x},\boldsymbol{\theta}^{J})$ are modeled using linear function approximators, that is, $\boldsymbol{g}(\boldsymbol{x},\boldsymbol{\theta}^{f})=\boldsymbol{\phi}_{f}(\boldsymbol{x})\boldsymbol{\theta}^{f}$ and $V(\boldsymbol{x},\boldsymbol{\theta}^{J})=\boldsymbol{\phi}_{J}(\boldsymbol{x})^{\prime}\boldsymbol{\theta}^{J}$, where $\boldsymbol{\phi}_{f}(\boldsymbol{x}):\mathbb{R}^{n}\rightarrow\mathbb{R}^{r\times p}$ and $\boldsymbol{\phi}_{J}(\boldsymbol{x}):\mathbb{R}^{n}\rightarrow\mathbb{R}^{s}$ are manually chosen basis functions satisfying $\parallel\boldsymbol{\phi}_{f}(\boldsymbol{x})\parallel\leq 1$ and $\parallel\boldsymbol{\phi}_{J}(\boldsymbol{x})\parallel\leq 1$.
Given that the proposed framework seeks to minimize (4) by simultaneously tuning $\boldsymbol{\theta}^{f}$, $\boldsymbol{\theta}^{J}$, and $\boldsymbol{\theta}^{\mu}$, where $\boldsymbol{\theta}^{f}$ and $\boldsymbol{\theta}^{J}$ are updated independently while $\boldsymbol{\theta}^{\mu}$ is updated based on $\boldsymbol{\theta}^{f}$ and $\boldsymbol{\theta}^{J}$, it is essential to first analyze the convergence properties of the parameters $\boldsymbol{\theta}_{k+1}^{f}$ and $\boldsymbol{\theta}_{k+1}^{J}$. Prior to presenting the main results regarding the approximated policy gradient, we establish the following key result related to the dynamics approximation:
@@
If Assumption 1 holds and $\boldsymbol{\theta}^{f}$ is updated following (12), for constant step size satisfying $0<\alpha_{f}<\frac{1}{2L_{f1}}$, one has $\lim_{k\rightarrow\infty}\parallel\boldsymbol{\theta}_{k}^{f}-\boldsymbol{\theta}^{f*}\parallel^{2}=\frac{\alpha_{f}L_{f2}}{2L_{f1}}$, where $L_{f1}$ is a constant, and $L_{f2}$ denotes a constant determined by the residual vectors associated with the least squares solutions in (13), as detailed in the Appendix. To obtain $\lim_{k\rightarrow\infty}\parallel\boldsymbol{\theta}_{k}^{f}-\boldsymbol{\theta}^{f*}\parallel^{2}=0$, one needs to appropriately choose a diminishing step size satisfying $0<\alpha_{k}^{f}<\frac{1}{2L_{f1}}$, with $\sum_{k=0}^{\infty}\alpha_{k}^{f}=\infty$ and $\sum_{k=0}^{\infty}(\alpha_{k}^{f})^{2}<\infty$.
The proof of Lemma 1 is provided in the Appendix. Based on Lemma 1, we establish the following corollary:
Corollary 1
Under the same assumptions as in Lemma 1, choosing the step size $\alpha_{k}^{f}=\frac{\beta_{f}}{2+k}$ with $\beta_{f}=\frac{1}{L_{f1}}$ yields
@@
where $\nu_{f}=\max\{\beta_{f}^{2}L_{f2},\quad 2\parallel\boldsymbol{\theta}_{0}^{f}-\boldsymbol{\theta}^{f*}\parallel^{2}\}$.
Remark 3
Since $L_{f2}$ in (18) is derived from the residual vectors from the least squares solutions in (13), it generally results in a smaller value of $L_{f2}$ compared to using the arbitrarily assigning the constant matrices $A$, $B$, and $C$. This reduction in $L_{f2}$ consequently leads to a tighter upper bound in (18), thereby expediting convergence of $\boldsymbol{\theta}_{k}^{f}\rightarrow\boldsymbol{\theta}^{f*}$.
As TD error learning is a well-established method, we refer to the seminal work [39] for a detailed theoretical analysis. In particular, Theorem $1$ in [39] proves that under certain assumptions, when $\boldsymbol{\theta}_{k}^{J}$ is updated according to (14), it holds that $\lim_{k\rightarrow\infty}\parallel\boldsymbol{\theta}_{k}^{J}-\boldsymbol{\theta}^{J*}\parallel^{2}=0$. A detailed explanation of this result is provided in the Appendix.
To examine the convergence behavior of $\boldsymbol{\theta}_{k}^{\mu}$, we first establish the following consequence to demonstrate the Lipschitz continuity of the introduced $\hat{J}_{\mu,t^{*}}$ in (17):
@@
If Assumption 1 holds, then the introduced $\hat{J}_{\mu,t^{*}}$ in (17) and its gradient $\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu,t^{*}}$ are also Lipschitz continuous with respect to $\boldsymbol{\theta}^{\mu}$, with Lipschitz constants $L_{J\theta}$ and $L_{J\theta\theta}$, respectively, as detailed in the Appendix.
The proof of Lemma 2 is provided in the Appendix. Building upon the preceding results, as $\boldsymbol{\theta}_{k}^{f}\rightarrow\boldsymbol{\theta}^{f*}$ and $\boldsymbol{\theta}_{k}^{J}\rightarrow\boldsymbol{\theta}^{J*}$, we have the following convergence properties of the approximated policy gradient:
@@
Under the same assumptions as in Lemma 1 and Lemma 2, if the parameters $\boldsymbol{\theta}_{k}^{f}$, $\boldsymbol{\theta}_{k}^{J}$, and $\boldsymbol{\theta}_{k}^{\mu}$ are updated according to (12), (14), and (15), respectively, using the step sizes specified in Corollary 1 and Lemma 3. To attain $\lim_{k\rightarrow\infty}\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})\parallel^{2}=0$, it is necessary to use diminishing step size $\alpha_{k}^{\mu}$ satisfying $\sum_{k=0}^{\infty}\alpha_{k}^{\mu}=\infty$ and $\sum_{k=0}^{\infty}(\alpha_{k}^{\mu})^{2}<\infty$.
The proof of Theorem 1 is provided in the Appendix. As a direct consequence of Theorem 1, we state the following corollary:
Corollary 2
If the assumptions in Theorem 1 hold, and the step sizes are chosen as $\alpha_{k}^{f}=\frac{\beta_{f}}{2+k}$ from Corollary 1, $\alpha_{J}=(1-\gamma)/4$ from Lemma 3, and $\alpha_{k}^{\mu}=(k+1)^{-1/4}$, then the estimated policy gradient in (16) has the following property:
$\min_{k\in\{0,1,2,\cdots,K-1\}}\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})\parallel^{2}$
@@
where $L_{b}=\frac{3}{4}(\hat{J}_{\mu}(\boldsymbol{\theta}_{0}^{f},\boldsymbol{\theta}_{0}^{J},\boldsymbol{\theta}_{0}^{\mu})-\hat{J}_{\mu}(\boldsymbol{\theta}^{f*},\boldsymbol{\theta}^{J*},\boldsymbol{\theta}^{\mu*})+C_{1}(\frac{2}{1-\sqrt{1-\nu_{J}}}+(1-\nu_{J})^{K/2})\parallel\boldsymbol{\theta}_{0}^{J}-\boldsymbol{\theta}^{J*}\parallel)$, and constants $L_{a}$, $L_{c}$ and $C_{1}$ are specified in the Appendix, while $10<\nu_{J}<$ 1 is a constant defined in Lemma 3.
Remark 4
Theorem 1 and Corollary 2 establish the convergence properties of the estimated policy gradient $\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}_{\mu}$ in (16) under appropriately chosen step sizes. It states that, under certain assumptions, if the dynamics and cost function are successfully learned, that is, $\boldsymbol{x}(t+1)=C^{*}(A^{*}\boldsymbol{g}(\boldsymbol{x}(t),\boldsymbol{\theta}^{f*})+B^{*}\boldsymbol{u}(t))$ and $J_{\mu,t^{*}}=V(\boldsymbol{x}(t^{*}),\boldsymbol{\theta}^{J*})$, then the proposed framework can achieve a globally optimal closed-loop controller, provided that the functions $c$ and $V$ are convex. In the case where these functions are non-convex, the framework is capable of obtaining a locally optimal closed-loop controller.
In the following section, we validate the theoretical results presented above and demonstrate the performance of the proposed framework through numerical simulations.
@@ 4.1 Online Implementation
We now present an online algorithm for the efficient implementation of the proposed policy gradient estimation framework, which employs the widely used mini-batch gradient descent method [40] to update the parameters $\boldsymbol{\theta}_{k}^{f}$, $\boldsymbol{\theta}_{k}^{J}$, and $\boldsymbol{\theta}_{k}^{\mu}$, and is detailed below.
Data generation. To ensure the sufficient exploration of the system state space $\mathcal{X}$, we adopt the widely-used method from [8], where an exploration policy is implemented by introducing noise $W(t)\in\mathcal{W}\subset\mathbb{R}^{m}$, sampled from a noise process, into the policy in (3) at each time step $t$ while updating $\boldsymbol{\theta}_{k}^{\mu}$ following (5). Common choices for $W(t)$ include the Ornstein-Uhlenbeck process [41] or modeling it as a Gaussian distribution. Specifically, let $t_{i}$ denote the system time at the $i$-th observation. Given system state $\boldsymbol{x}(t_{i})=\boldsymbol{x}_{t_{i}}$, by applying
@@
For $i>N$, the earliest $i-N$ observed tuples are discarded to accommodate new data. The value of $N$ is typically chosen to be large. At any iteration $k$, a data batch of $\tilde{N}$ tuples ($\tilde{N}\ll N$) is sampled from $\mathcal{D}_{i}$, with its index set $\mathcal{I}_{k}$ drawn uniformly from all subsets of $\{1,2,\cdots,i\}$ with size $\tilde{N}$, provided that $i\geq\tilde{N}$. This batch is then used to estimate the gradients in (12), (14), and (16). Note that, under the online setting, data collection and parameter updates are performed concurrently.
Parameters tuning using sampled data batches. Given the above sampled data batch from $\mathcal{D}_{i}$ at any iteration $k$, we construct the following data matrices, which are exclusively utilized within the scope of this subsection:
${\mathbf{X}}_{k}$ $=[\boldsymbol{x}_{1},\boldsymbol{x}_{2},\cdots,\boldsymbol{x}_{\tilde{N}}]\in\mathbb{R}^{n\times\tilde{N}},$ (20)
$\mathbf{U}_{k}$ $=[\boldsymbol{u}_{1},\boldsymbol{u}_{2},\cdots,\boldsymbol{u}_{\tilde{N}}]\in\mathbb{R}^{m\times\tilde{N}},$
$\bar{{\mathbf{X}}}_{k}$ $=[\boldsymbol{x}_{1}^{+},\boldsymbol{x}_{2}^{+},\cdots,\boldsymbol{x}_{\tilde{N}}^{+}]\in\mathbb{R}^{n\times\tilde{N}},$
${\mathbf{G}}_{k}$ $=[\boldsymbol{g}(\boldsymbol{x}_{1},\boldsymbol{\theta}_{k}^{f}),\boldsymbol{g}(\boldsymbol{x}_{2},\boldsymbol{\theta}_{k}^{f}),\cdots,\boldsymbol{g}(\boldsymbol{x}_{\tilde{N}},\boldsymbol{\theta}_{k}^{f})]\in\mathbb{R}^{r\times\tilde{N}},$
$\bar{{\mathbf{G}}}_{k}$ $=[\boldsymbol{g}(\boldsymbol{x}_{1}^{+},\boldsymbol{\theta}_{k}^{f}),\boldsymbol{g}(\boldsymbol{x}_{2}^{+},\boldsymbol{\theta}_{k}^{f}),\cdots,\boldsymbol{g}(\boldsymbol{x}_{\tilde{N}}^{+},\boldsymbol{\theta}_{k}^{f})]\in\mathbb{R}^{r\times\tilde{N}},$
where $(\boldsymbol{x}_{\tilde{N}},\boldsymbol{u}_{\tilde{N}},\boldsymbol{x}_{\tilde{N}}^{+})$ represents the $\tilde{N}$-th data tuple in the sampled batch, as used exclusively in (20) for notational convenience. This notation is employed because the objective is to find the mapping between $\boldsymbol{x}_{t},\boldsymbol{u}_{t}$ and $\boldsymbol{x}_{t}^{+}$ without regard to the specific time at which it is recorded.
If the matrices $\bar{{\mathbf{G}}}_{k}\in\mathbb{R}^{r\times\tilde{N}}$ and $\begin{bmatrix}{\mathbf{G}}_{k}\\ \mathbf{U}_{k}\end{bmatrix}\in\mathbb{R}^{(r+m)\times\tilde{N}}$ are with full row rank, then the parameter $\boldsymbol{\theta}_{k}^{f}$ is updated in the same manner as (12) with its gradient computed using the sampled data batch, where $[A_{k},B_{k}]$ and $C_{k}$ are computed according to (13) using the data matrices defined in (20).
After updating $\boldsymbol{\theta}_{k}^{f}$, the parameters $\boldsymbol{\theta}_{k}^{J}$ and $\boldsymbol{\theta}_{k}^{\mu}$ are updated in a manner analogous to (14) and (15), respectively, using gradients computed over the sampled data batch. In summary, we introduce an online PGDK in Algorithm 2.
1 Initialize $\boldsymbol{\theta}_{0}^{f}\in\mathbb{R}^{p}$, $\boldsymbol{\theta}_{0}^{J}\in\mathbb{R}^{s}$, and $\boldsymbol{\theta}_{0}^{\mu}\in\mathbb{R}^{q}$ for $\boldsymbol{g}(\cdot,\boldsymbol{\theta}^{f})$, $V(\cdot,\boldsymbol{\theta}^{J})$, and $\boldsymbol{\mu}(\cdot,\boldsymbol{\theta}^{\mu})$, respectively. Set the iteration index $k=0$ and step size sequences $\{\alpha_{k}^{f}\}_{k=0}^{K}$, $\{\alpha_{k}^{J}\}_{k=0}^{K}$, $\{\alpha_{k}^{\mu}\}_{k=0}^{K}$. Specify discount factor $\gamma$, number of episodes $E$, task horizon $T$, batch size $\tilde{N}$, time-decay function $\sigma(t)$, and data memory.
@@
6             Sample $\tilde{N}$ data tuples uniformly from the data memory.
7             Update $\boldsymbol{\theta}_{k}^{f}$ analogous to (12) using the batch gradient, where $A_{k}$, $B_{k}$, and $C_{k}$ are computed following (13) using (20).
8             Update $\boldsymbol{\theta}_{k}^{J}$ and $\boldsymbol{\theta}_{k}^{\mu}$ in a manner analogous to (14) and (15), respectively, using the gradient computed over the sampled data batch.
@@
Algorithm 2 Online policy gradient based on deep Koopman representation (Online-PGDK)

### v3

@@ 3.3 Analysis
In this subsection, we analyze the convergence of the proposed framework in (15) for a given dataset $\mathcal{D}$. To this end, we first introduce the following assumption and definition:
@@
The unknown dynamics $\boldsymbol{f}$ in (1) can be rewritten as a deep Koopman dynamics in (7), and $J_{t^{*}}$ in (2) can be expressed by the parameterized function $V(\boldsymbol{x}_{t^{*}},\boldsymbol{\theta}^{J*})$.
As noted in [brunton2016koopman], a finite-dimensional Koopman representation in (8) is achievable only if $\boldsymbol{f}$ has a single isolated fixed point.
Definition 1
A sequence of step sizes $\{\alpha_{k}\}_{k=0}^{\infty}$ with $\alpha_{k}\geq 0$ is said to satisfy the Robbins–Monro (RM) condition if $\sum_{k=0}^{\infty}\alpha_{k}=\infty$ and $\sum_{k=0}^{\infty}\alpha_{k}^{2}<\infty$.
Since $\boldsymbol{\theta}^{\mu}$ is updated based on $\boldsymbol{\theta}^{f}$ and $\boldsymbol{\theta}^{J}$, it is first necessary to analyze the convergence of $\boldsymbol{\theta}_{k}^{f}$ in (12) and $\boldsymbol{\theta}_{k}^{J}$ in (14). Thus, we establish the following key result regarding the dynamics approximation with respect to $\boldsymbol{\theta}^{f}$:
@@
If Assumption 1 holds and $\boldsymbol{\theta}_{k}^{f}$ is updated following (12) with step size $\alpha_{k}^{f}=\frac{1}{L_{f1}(2+k)}$, then
@@
where $\nu_{f}=\max\{L_{f2}/L_{f1}^{2},2\parallel\boldsymbol{\theta}_{0}^{f}-\boldsymbol{\theta}^{f*}\parallel^{2}\}$, $L_{f1}$ is a constant, and $L_{f2}$ is a constant determined by the residual vectors from the least squares solutions in (13). Furthermore, to guarantee $\lim_{k\rightarrow\infty}\parallel\boldsymbol{\theta}_{k}^{f}-\boldsymbol{\theta}^{f*}\parallel^{2}=0$, the step size sequence $\{\alpha_{k}^{f}\}_{k=0}^{\infty}$ must satisfy the RM condition.
The proof of Lemma 1 is provided in the Appendix. Lemma 1 says that, under certain assumptions, the update rule in (12) converges sublinearly. Following established TD-error results, we summarize from [bhandari2018finite]:
@@
[Theorem 11, [bhandari2018finite]] If the dynamics (1) following the policy $\boldsymbol{\mu}$ is ergodic with a unique stationary distribution $\pi(\boldsymbol{x}):\mathbb{R}^{n}\rightarrow\mathbb{R}$. That is, for any two states $\boldsymbol{x}_{1},\boldsymbol{x}_{2}:\pi(\boldsymbol{x}_{2})=\lim_{t\rightarrow\infty}\mathbb{P}(\boldsymbol{x}_{t}=\boldsymbol{x}_{2}|\boldsymbol{x}_{0}=\boldsymbol{x}_{1})$. Let $\Sigma=\sum_{\boldsymbol{x}\in\mathcal{X}}\pi(\boldsymbol{x})\boldsymbol{\phi}_{J}(\boldsymbol{x})\boldsymbol{\phi}_{J}(\boldsymbol{x})^{\prime}$, and denote $\omega_{J}$ as the smallest eigenvalue of $\Sigma$, and $\nu_{J}=(1-\gamma)^{2}\omega_{J}/4$. Given $\omega_{J}>0$ and a constant step size $\alpha_{J}=(1-\gamma)/4$, following (14) yields:
Building on the preceding results, the approximated policy gradient satisfies the following convergence properties:
@@
If Assumption 1 holds and the parameters $\boldsymbol{\theta}_{k}^{f}$, $\boldsymbol{\theta}_{k}^{J}$, and $\boldsymbol{\theta}_{k}^{\mu}$ are updated according to (12), (14), and (15) with step sizes $\alpha_{k}^{f}=\frac{1}{L_{f1}(2+k)}$, $\alpha_{J}=(1-\gamma)/4$, and $\alpha_{k}^{\mu}=(k+1)^{-1/4}$, respectively. Then, the policy gradient satisfies
$\min_{k\in\{0,1,2,\cdots,K-1\}}\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})\parallel^{2}$
@@
where $L_{a}$, $\lim_{k\rightarrow\infty}L_{b}$, and $L_{c}$ are constants defined in the Appendix. The constant $L_{f1}$ is as specified in Lemma 1. Moreover, to ensure $\lim_{k\rightarrow\infty}\parallel\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}(\boldsymbol{\theta}_{k+1}^{f},\boldsymbol{\theta}_{k+1}^{J},\boldsymbol{\theta}_{k}^{\mu})\parallel^{2}=0$, $\{\alpha_{k}^{\mu}\}_{k=0}^{\infty}$ needs to satisfy the RM condition.
The proof of Theorem 1 is provided in the Appendix. Theorem 1 establishes that the policy gradient converges when the policy is updated on a faster time scale than the DKO learner. Furthermore, it shows that the proposed method achieves a globally optimal policy provided that the functions $c$, $\boldsymbol{g}$, and $V$ are convex. In the non-convex case, the method converges only to a locally optimal policy.
If Assumption 1 does not hold, then $\nabla_{\boldsymbol{\theta}^{\mu}}\hat{J}$ in (16) only approximates $\nabla_{\boldsymbol{\theta}^{\mu}}J_{t^{*}}$ in (6) with a deterministic error $\epsilon_{k}$. Since $\boldsymbol{f}$, $J_{t^{*}}$, $\boldsymbol{g}$, and $V$ are Lipschitz continuous, there exists $\epsilon>0$ such that $\parallel\epsilon_{k}\parallel\leq\epsilon$. According to [bertsekas2000gradient], if $J_{t^{*}}$ is strongly convex and a time-decaying step size $\alpha_{k}^{\mu}\rightarrow 0$ is employed, then $\lim_{k\rightarrow\infty}\parallel\boldsymbol{\theta}_{k}^{\mu}-\boldsymbol{\theta}^{\mu*}\parallel^{2}$ converges to a positive constant that is proportional to $\epsilon^{2}$.
@@ 4.1 Online Implementation
In practical scenarios, the offline approach presents challenges, particularly when $N$ is large, making dataset acquisition and gradient computation computationally intensive. Moreover, discrepancies between the current policy and the behavior policy employed during offline data collection may impede the identification of the closed-loop optimal controller, as noted in [agarwal2020optimistic]. To address these limitations arising from training data constraints and to enhance implementation efficiency, we introduce an online variant of PGDK that extends Algorithm 1. In this formulation, instead of utilizing a fixed offline dataset $\mathcal{D}$ to estimate the policy gradient, the proposed online method employs the mini-batch gradient descent scheme [mnih2013playing], widely employed in RL. Specifically, at each iteration $k$, the policy parameters are updated using gradients computed from a sampled mini-batch, as described below.
Data batches sampling. To ensure the sufficient exploration of the system state space $\mathcal{X}$, we adopt the widely-used method from [lillicrap2015continuous], where an exploration policy is implemented by introducing noise $W(t)\in\mathcal{W}\subset\mathbb{R}^{m}$, sampled from a noise process, into the policy in (3) at each time step $t$ while updating $\boldsymbol{\theta}_{k}^{\mu}$ following (5). Common choices for $W(t)$ include the Ornstein-Uhlenbeck process [uhlenbeck1930theory] or modeling it as a Gaussian distribution. Specifically, let $t_{i}$ denote the system time at the $i$-th observation. Given system state $\boldsymbol{x}(t_{i})=\boldsymbol{x}_{t_{i}}$, by applying
@@
For $i>N$, the earliest iNi-N observed tuples are discarded to accommodate new data. The value of $N$ is typically chosen to be large. At any iteration $k$, a mini-batch $\mathcal{D}_{k}$ of $\tilde{N}$ tuples ($\tilde{N}\ll N$) is sampled from $\mathcal{D}_{i}$, with its index set $\mathcal{I}_{k}$ drawn uniformly from all subsets of $\{1,2,\cdots,i\}$ with size $\tilde{N}$, provided that $i\geq\tilde{N}$. This sampled $\mathcal{D}_{k}$ is then used to construct (11) and estimate the gradients in (12), (14), and (16).
Under the online setting, data collection and parameter updates are performed concurrently. Moreover, the gradients of (12), (14), and (16) computed over $\mathcal{D}_{k}$ serve as an unbiased estimation of those obtained from the full dataset $\mathcal{D}_{i}$.
In summary, Algorithm 2 presents the online variant of PGDK, which implements the PGDK framework by computing gradients over sampled mini-batches rather than a fixed offline dataset. The overall structure of the algorithm is illustrated in Fig. 2.
Figure 2: PGDK framework under online implementation.
1Initialize $\boldsymbol{\theta}_{0}^{f}\in\mathbb{R}^{p}$, $\boldsymbol{\theta}_{0}^{J}\in\mathbb{R}^{s}$, and $\boldsymbol{\theta}_{0}^{\mu}\in\mathbb{R}^{q}$ for $\boldsymbol{g}(\cdot,\boldsymbol{\theta}^{f})$, $V(\cdot,\boldsymbol{\theta}^{J})$, and $\boldsymbol{\mu}(\cdot,\boldsymbol{\theta}^{\mu})$, respectively. Set the iteration index $k=0$ and step size sequences $\{\alpha_{k}^{f}\}_{k=0}^{K}$, $\{\alpha_{k}^{J}\}_{k=0}^{K}$, $\{\alpha_{k}^{\mu}\}_{k=0}^{K}$. Specify discount factor $\gamma$, number of episodes $E$, task horizon $T$, batch size $\tilde{N}$, time-decay function $\sigma(t)$, and data memory.
@@
6      Sample mini-batch $\mathcal{D}_{k}$ of $\tilde{N}$ tuples uniformly from the data memory.
7      Update $\boldsymbol{\theta}_{k}^{f}$ analogous to (12) using the batch gradient, where $A_{k}$, $B_{k}$, and $C_{k}$ are computed following (13) using $\mathcal{D}_{k}$.
8      Update $\boldsymbol{\theta}_{k}^{J}$ and $\boldsymbol{\theta}_{k}^{\mu}$ in a manner analogous to (14) and (15), respectively, using the gradient computed over $\mathcal{D}_{k}$.
@@
Algorithm 2 Policy Gradient with Deep Koopman Representation (PGDK) — online implementation